# SynthVLM: Towards High-Quality and Efficient Synthesis of Image-Caption Datasets for Vision-Language Models

### v4

@@ 3.3. High Quality Synthetic Dataset
In this section, we compare commonly used image-caption datasets with the SynthVLM-100K dataset. The synthetic data offers high image quality, excellent image-text alignment, superior model ratings, and robust data privacy protection.
High Image Quality
As illustrated in Figure 2, SynthVLM markedly improves image quality by generating images at a resolution of 1024x1024 pixels. This high resolution addresses the prevalent issue of suboptimal image quality in existing datasets, providing high-quality image-caption pairs that are invaluable for training VLLMs. Additionally, SynthVLM-100K effectively eliminates artifacts such as watermarks and advertisements.
Excellent Image-Text Alignment.
Excellent Model Rating.
Since our data will be used for VLLMs training, we use VLLMs to evaluate the data quality. We selected 1K image-caption pairs and submitted the caption along with the synthetic image and the original image. We used human annotators, GPT-4 Vision and Intern-VL2 as the judge model and requested it to select the pair that exhibited higher alignment. The specific prompt used for this evaluation is illustrated in Figure 3. The results, presented in Table 2, demonstrate that images generated have better alignment with the caption.
@@
Protect Data Privacy.
A notable advantage of our dataset is its exclusive reliance on captions, which mitigates data privacy concerns. By not using existing images, we ensure that no sensitive or private information associated with those images is compromised. This approach adheres to data privacy best practices, ensuring that our dataset maintains both high quality and ethical integrity.
Through these two training stages, we successfully developed the VLLM model trained by SynthVLM-100K data. Our method is efficient, utilizing only 100K pre-training data while also protecting privacy by leveraging synthetic data. Additionally, SynthVLM provides a new paradigm for effective alignment between modalities in Vision Language Models using pure synthetic data.
@@ 4. Experiments
In this section, we utilize the image data synthesis system for various tasks. We then aim to answer the following questions to verify the effectiveness, efficiency, and privacy protection of our proposed SynthVLM: Q1: Can our SynthVLM-7B and SynthVLM-13B achieve SOTA performance compared to previous SOTA methods? Q2: How does the efficiency of our SynthVLM compare to previous methods? Q3: Can SynthVLM protect privacy while achieving state-of-the-art performance? Q4: Would using generated data in full impact the application of the model in real-world scenarios? Q5: Do we need the generate module and the quality selection module to enhance model performance?
@@ 4.1. Experimental Settings
We use the LLaVA 1.5 model to validate the effectiveness of our dataset. For the Vision Encoder, we select CLIP 336px, and for the LLM, we use Vicuna 7B and Vicuna 13B. Our training parameters are consistent with LLaVA, and we adopt a two-stage training approach. In the pre-training stage, we train the projector to align the image and text modalities. The SynthVLM-100K dataset described in Section 3.3 is used for this purpose. In the SFT stage, we further train the projector alongside the LLM to improve visual understanding capabilities.
@@
For LLaVA-BaseLine-7B and LLaVA-BaseLine-13B, we use the LLaVA 558k dataset for pre-training and the LLaVA 665k dataset for SFT. For SynthVLM-7B and SynthVLM-13B, we use the SynthVLM-100K dataset for pre-training and the LLaVA 665k dataset for SFT.
@@
We select benchmarks for both visual understanding and language understanding. For visual understanding, we choose SQAI, MMVet, VizWiz, VQAv2, GQA, MME, and PoPE for a comprehensive evaluation. For language benchmarks, we select MMLU and SQA to assess language understanding abilities.
@@ 4.2. Synthetic Data Achieves SOTA Performance
To address Q1, we trained the model described in experimental settings. From Table 5, it is evident that SynthVLM model outperforms the Baseline across all evaluation benchmarks on both 7B and 13B. Specifically, SynthVLM model achieves SOTA results on visual benchmarks such as SQAI, MMVet, VizWiz, VQAv2, GQA, MME, MMB, and PoPE. Furthermore, SynthVLM model also excels in pure language benchmarks, demonstrating superior performance in SQA and MMLU, thus showcasing its comprehensive capabilities in both vision and language tasks.
@@ 4.3. Efficient Vision Language Alignment
Computational Resource Usage
By integrating a data selection module, our approach utilizes only 19% of the LLAVA data and 10% of the original synthetic data while achieving SOTA performance. This demonstrates that our data selection method can reduce computational usage by more than 80%.
Data Utilization Efficiency
Table 4 compares the data utilization of our proposed method with the conventional BLIP-LCS method, which employs the BLIP model for caption generation. Our approach is significantly more efficient, requiring only 330 MB for captions to generate 1,000,000 image-caption pairs, compared to the traditional methods, which may exceed 50 GB for images. This substantial difference highlights the efficiency of relying solely on captions rather than images in generating image-caption pairs.
Models LLM SQA SQAI MMVet VizWiz VQAv2 GQA MMB MMEP MMEC PoPE MMLU
w/o generation Vicuna-1.5-7B 69.3$\downarrow$ 67.0$\downarrow$ 31.2$\downarrow$ 46.8$\downarrow$ 79.3$\downarrow$ 62.9$\downarrow$ 66.2$\downarrow$ 1488.8$\downarrow$ 327.5$\downarrow$ 86.2$\downarrow$ 39.1$\downarrow$
w/o selection Vicuna-1.5-7B 69.9$\downarrow$ 67.7$\downarrow$ 30.2$\downarrow$ 50.2 79.1$\downarrow$ 62.2$\downarrow$ 63.5$\downarrow$ 1421.7$\downarrow$ 301.8$\downarrow$ 87.3 40.6$\downarrow$
w/o generation Vicuna-1.5-13B 73.6$\downarrow$ 71.4$\downarrow$ 33.0$\downarrow$ 53.6$\downarrow$ 80.0 63.4$\downarrow$ 67.5$\downarrow$ 1514.3$\downarrow$ 295.7$\downarrow$ 88.2$\downarrow$ 53.6$\downarrow$
w/o selection Vicuna-1.5-13B 74.1$\downarrow$ 70.5$\downarrow$ 35.6 53.2$\downarrow$ 79.7$\downarrow$ 63.1$\downarrow$ 67.5$\downarrow$ 1512.7$\downarrow$ 303.2$\downarrow$ 86.9$\downarrow$ 53.0$\downarrow$
Models LLM SQA MMLU
Avg STEM Humanities Social Sciences Other
SynthVLM-7B Vicuna-1.5-7B 70.4 41.2 31.7 37.4 47.0 50.2
w/o generation Vicuna-1.5-7B 69.3$\downarrow$ 39.1$\downarrow$ 30.0$\downarrow$ 36.6$\downarrow$ 43.1$\downarrow$ 47.3$\downarrow$
w/o selection Vicuna-1.5-7B 69.9$\downarrow$ 40.6$\downarrow$ 30.8$\downarrow$ 37.2$\downarrow$ 45.3$\downarrow$ 48.9$\downarrow$
SynthVLM-13B Vicuna-1.5-13B 74.9 54.6 45.0 49.3 64.0 62.2
w/o generation Vicuna-1.5-13B 74.1$\downarrow$ 53.6$\downarrow$ 43.5$\downarrow$ 48.2$\downarrow$ 63.1$\downarrow$ 61.8$\downarrow$
w/o selection Vicuna-1.5-13B 73.6$\downarrow$ 53.0$\downarrow$ 42.9$\downarrow$ 46.8$\downarrow$ 63.8$\downarrow$ 61.3$\downarrow$
4.4. Privacy Protection Synthetic Data
To address Q3, we compare the synthetic image and the original image in Figure 4. Synthetic data offers significant advantages in protecting data privacy. In Figure 4, synthetic images in (a) show vehicles and tickets without revealing real license plates and ticket information, ensuring privacy protection. In contrast, original images in (b) display actual license plates and ticket information, which can potentially lead to privacy issues.
@@ 4.5. Capability to Solve Real-World Problems
A critical issue when using generated data is whether the model loses its ability to solve real-world problems. To address Q4, we utilized the MME Benchmark to assess the model’s performance in various real-world scenarios, such as artwork, celebrity, code reasoning, landmarks, and posters. The specific results are presented in Table 8. Using SynthVLM-100K, our model performs consistently well, matching or even surpassing the baseline across most real-world tasks. We attribute this success to the diversity of the captions and the generalization ability of the Diffusion model.
@@ 4.6. Ablation Study
To address Q5, we conducted the following ablation study. Specifically, we conducted an ablation study where we removed the data generation module and the data selection module separately to evaluate their individual contributions to the effectiveness of our data generation pipeline. In this section, we controlled the number of samples in each experimental group to be 100K.
@@
The exclusion of the data generation module significantly impacts the model’s performance, as illustrated in Tables 8 and 9, labeled as ”w/o generation”. The variant without this module demonstrates markedly lower accuracy across all benchmarks. These results emphasize the crucial role of the data generation process in sustaining the high performance of the SynthVLM model. This also underscores SynthVLM’s potential in constructing highly aligned datasets.
@@
Overall, the ablation study highlights the critical role of the data generation and the data selection modules in SynthVLM. These experiments provide valuable insights into the contributions of each module, guiding future improvements and optimizations of the SynthVLM model.
@@ 5. Conclusion
Data is crucial for training VLLMs. However, high-quality web data is limited. In this paper, we introduce a novel image synthesis pipeline to generate pre-training data for VLLMs. Our synthetic images are of higher quality as they are free from watermarks and advertisements. Furthermore, our synthetic image-caption pairs exhibit superior CLIPScore and SSIMScore, reflecting better image-caption alignment and image quality. Remarkably, the SynthVLM model, trained with only 18% synthetic data, surpasses the LLaVA baseline model trained on the full dataset. Moreover, it efficiently achieves SOTA alignment capabilities while preserving the language proficiency of VLLMs. By employing synthetic datasets, SynthVLM also safeguards data privacy, underscoring the practicality and effectiveness of our data synthesis pipeline.

### v5

@@ 3.3. High Quality Synthetic Dataset
In this section, we compare commonly used image-caption datasets with the SynthVLM-100K dataset. The synthetic data offers high image quality, excellent image-text alignment and superior model ratings.
High Image Quality
As illustrated in Figure 1, SynthVLM markedly improves image quality by generating images at a resolution of 1024x1024 pixels. This high resolution addresses the prevalent issue of suboptimal image quality in existing datasets, providing high-quality image-caption pairs that are invaluable for training VLMs. Additionally, SynthVLM-100K effectively eliminates artifacts such as watermarks and advertisements.
Excellent Image-Text Alignment.
Excellent Model Rating.
Since our data will be used for VLMs training, we use VLMs to evaluate the data quality. We selected 1K image-caption pairs and submitted the caption along with the synthetic image and the original image. We used human annotators, GPT-4 Vision(OpenAI et al., 2024) and Intern-VL2(Wang et al., 2025) as the judge model and requested it to select the pair that exhibited higher alignment. The specific prompt used for this evaluation is illustrated in Figure 3. The results, presented in Table 6, demonstrate that images generated have better alignment with the caption.
Sample Model Sythetic Image win Web Image win
1K GPT4-Vision 633 367
1K InternVL2 692 308
1K Human 758 242
Through these two stages, we successfully developed SynthVLM-100K data. Our method is efficient, utilizing only 100K pre-training data. Additionally, SynthVLM provides a new paradigm for effective alignment between modalities in Vision Language Models using pure synthetic data.
Methods SynthVLM LLaVA w/o selection
Dataset Number (k) 100 558 1000
Data Usage 33MB 27GB 310MB
@@ 4. Experiments
In this section, we utilize the image data synthesis system for various tasks. We then aim to answer the following questions to verify the effectiveness and efficiency of our proposed SynthVLM: Q1: Can our SynthVLM-7B and SynthVLM-13B achieve SOTA performance compared to previous SOTA methods? Q2: How does the efficiency of our SynthVLM compare to previous methods? Q3: Would using generated data in full impact the application of the model in real-world scenarios? Q4: Do we need the generate module and the quality selection module to enhance model performance?
@@ 4.1. Experimental Settings
We use the LLaVA 1.5(Liu et al., 2023a) model to validate the effectiveness of our dataset. For the Vision Encoder, we select CLIP 336px, and for the LLM, we use Vicuna 7B and Vicuna 13B(Chiang et al., 2023). Our training parameters are consistent with LLaVA, and we adopt a two-stage training approach. In the pre-training stage, we train the projector to align the image and text modalities. The SynthVLM-100K dataset described in Section 3.3 is used for this purpose. In the SFT stage, we further train the projector alongside the LLM to improve visual understanding capabilities.
@@
For LLaVA-BaseLine, we use the LLaVA 558k dataset for pre-training and the LLaVA 665k dataset for SFT. For SynthVLM-7B and SynthVLM-13B, we use the SynthVLM-100K dataset for pre-training and the LLaVA 665k dataset for SFT.
@@
We select benchmarks for both visual understanding and language understanding. For visual understanding, we choose SQAI(Lu et al., 2022a), MMVet(Yu et al., 2024), VizWiz(Gurari et al., 2018), VQAv2(Shen et al., 2023), GQA(Hudson and Manning, 2019), MME(Fu et al., 2024), and PoPE(Li et al., 2023b) for a comprehensive evaluation. For language benchmarks, we select MMLU(Hendrycks et al., 2021) and SQA(Lu et al., 2022a) to assess language understanding abilities.
Models LLM SQA SQAI MMVet VizWiz VQAv2 GQA MMB MMEP MMEC PoPE MMLU
[…]
Models SQA MMLU
Avg STEM Humanities Social Sciences Other
SynthVLM-7B 70.4 41.2 31.7 37.4 47.0 50.2
[…]
SynthVLM-13B 74.9 54.6 45.0 49.3 64.0 62.2
[…]
@@ 4.2. Synthetic Data Achieves SOTA Performance
To address Q1, we trained the model described in experimental settings. From Table 2, it is evident that SynthVLM model outperforms the Baseline across all evaluation benchmarks on both 7B and 13B. SynthVLM model also excels in language benchmarks, demonstrating superior performance in SQA and MMLU, thus showcasing its comprehensive capabilities in both vision and language tasks.
@@ 4.3. Efficient Vision Language Alignment
As shown in Figure  7, by integrating a data selection module, our approach utilizes only 19% of the LLAVA data and 10% of the original synthetic data while achieving SOTA performance. This demonstrates that our data selection method can reduce computational usage by more than 80%.
@@ 4.4. Capability to Solve Real-World Problems
A critical issue when using generated data is whether the model loses its ability to solve real-world problems. To address Q3, we utilized the MME Benchmark to assess the model’s performance in various real-world scenarios, such as artwork, celebrity, code reasoning, landmarks, and posters. The specific results are presented in Table 5. Using SynthVLM-100K, our model performs consistently well, matching or even surpassing the baseline across most real-world tasks. We attribute this success to the diversity of the captions and the generalization ability of the Diffusion model.
@@ 4.5. Ablation Study
To address Q4, We conducted an ablation study where we removed the data generation module and the data selection module separately to evaluate their individual contributions to the effectiveness of our data generation pipeline. In this section, we controlled the number of samples in each experimental group to be 100K.
@@
The exclusion of the data generation module significantly impacts the model’s performance, as illustrated in Tables 8 and 9, labeled as ”w/o generation”. The variant without this module demonstrates lower accuracy, emphasizing the crucial role of the data generation process in sustaining the high performance of the SynthVLM model. This also underscores SynthVLM’s potential in constructing highly aligned datasets.
@@
Overall, the ablation study highlights the critical role of data generation and data selection in SynthVLM, providing valuable insights into the contributions of each module.
@@ 5. Conclusion
We propose SynthVLM, a novel pipeline for generating high-quality pre-training data for VLMs. Unlike existing datasets, our synthetic images are free from watermarks and advertisements, leading to cleaner visual inputs. The generated data demonstrate superior alignment and visual fidelity. Notably, the SynthVLM model—trained with only 18% synthetic data—outperforms the LLaVA baseline trained on the full dataset. It not only achieves state-of-the-art alignment performance but also preserves the language understanding capabilities of VLMs. Ablation studies further confirm the effectiveness of both the image generation and data selection modules, highlighting the practical value of our data synthesis approach.
Acknowledgments
This work is supported by the National Key R&D Program of China (2024YFA1014003), National Natural Science Foundation of China (92470121, 62402016), and High-performance Computing Platform of Peking University.
Appendix
Appendix A Preliminary
A.1. Diffusion Model
$q(x_{1},\ldots,x_{T}\mid x_{0}):=\prod_{t=1}^{T}q(x_{t}\mid x_{t-1}),$
where $\beta_{t}$ controls the variance of the noise added at each step.
The distribution after $t$ steps can be written as:
where $\bar{\alpha}_{t}=\prod_{i=1}^{t}(1-\beta_{i})$.
While DDPMs have shown promising results, several improvements have been proposed to enhance their efficiency  (Song et al., 2021; Zhu et al., 2024) and sample quality  (Nichol and Dhariwal, 2021; Dhariwal and Nichol, 2021). The superior performance of diffusion models has been leveraged in various sub-tasks, including image generation, image translation, inpainting  (Lu et al., 2022b; Su et al., 2023).
A.2. Vision Language Models
The integration of visual knowledge into large language models (LLMs) has become a pivotal area of research due to the rapid advancements in LLMs. VLMs combine vision information from vision encoders with LLMs, thus enabling these models to process and interpret visual inputs for various visual tasks (Liu et al., 2023c; Zhang et al., 2022; Li et al., 2022b) with enhanced accuracy and efficiency. Pioneering frameworks like CLIP (Radford et al., 2021) leverage contrastive learning on expansive image-caption datasets to align modalities, forming the groundwork for cross-modal comprehension. Various adapters (Liu et al., 2023b, a; Li et al., 2023c, 2022a; Jian et al., 2023; Lu et al., 2023) are introduced to further integrate different modalities. For example, LLaVA (Liu et al., 2023b, a) employs a straightforward MLP to inject the vision information into LLMs. Whereas more complex implementations like the Q-Former in BLIP (Li et al., 2022a, 2023c) utilize cross-attention to enhance modality integration.
Recent studies  (Wang et al., 2024b; Chen et al., 2023b; Liu et al., 2023b, a; Li et al., 2023g) aims to boost VLM performance by focusing on the quality of both pre-training and fine-tuning datasets. Models like LLaVA  (Liu et al., 2023b, a) and ShareGPT4V  (Chen et al., 2023b) have shown remarkable advancements in understanding and following complex instructions through instruction tuning. Although these improvements help align the vision modality and establish a solid basis for cross-modal comprehension, they require extensive datasets for training and could potentially diminish the model’s language capabilities.
Appendix B Implementation Details
B.1. Data Generation
In this section, we detail the hyperparameters and procedures used for data generation.
We employed the Stable Diffusion XL(SDXL) model for image synthesis, following the framework outlined by the original authors (Podell et al., 2023). To identify the optimal parameter configuration for our use case, we conducted a grid search strategy aimed at maximizing the CLIPScore for evaluating the semantic alignment between generated images and their corresponding textual descriptions.
Specifically, we randomly sampled 1k captions from our caption pool and used these samples to systematically evaluate different combinations of generation parameters. The grid search allowed us to empirically determine the most effective configuration for producing high-quality, semantically relevant synthetic images.
Based on this optimization process, we configured SDXL with 60 sampling steps. All images were generated at a resolution of 1024×1024 pixels. These configurations consistently yielded superior quality.
B.2. Data Selection
Table 10. Metric and Prompt used for Caption Filtering
Caption Filtering ## Rule-Based Metrics
Alphanumeric Filter: Tokenization: false, Min ratio: 0.60 Character Repetition Filter: Rep length: 10, Max ratio: 0.09373663 Flagged Words Filter: Language: en, Tokenization: false, Max ratio: 0.0 Perplexity Filter: Language: en, Max perplexity: 5500.0 Special Characters Filter: Min ratio: 0.16534802, Max ratio: 0.42023757 Word Repetition Filter: Language: en, Tokenization: false, Rep length: 10, Max ratio: 0.03085751 Image-Text Matching Filter: HF BLIP: Salesforce/blip-itm-base-coco, Min score: 0.8, Max score: 1.0, Horizontal flip: false, Vertical flip: false, Reduce mode: avg, Any or all: any, Mem required: 1500MB Image-Text Similarity Filter: HF CLIP: openai/clip-vit-base-patch32, Min score: 0.28 ## Prompt
Assume you are an expert in the field of AI image generation. Your goal is to select high-descriptive prompts that will enable the successful generation of images. I will provide you with a specific descriptive prompt, and your task is to evaluate it thoroughly. Consider the prompt’s level of detail, its logical coherence, and the clarity with which it describes the desired image. It is essential to assess whether the prompt contains sufficient information to guide the diffusion model effectively, ensuring that it can produce an image that meets expectations. You should only respond with Yes or No.
In this section, we describe the strategy and prompts used for data selection. Our goal was to curate a high-quality dataset that aligns closely with our generation objectives.
To achieve this, we employed a two-stage filtering process combining heuristic rules and large language model (LLM)-based evaluation. The specific filtering rules and prompt templates are detailed in Table 10.
For heuristic filtering, we utilized the Data-Juicer framework, which offers a modular and scalable pipeline for rule-based data preprocessing. This allowed us to implement filters targeting criteria such as minimum caption length, syntactic completeness, and lexical diversity. Additionally, we removed low-information and repetitive entries to enhance the overall quality of the dataset.
Following this, we conducted LLM-based filtering using LLaMA3-70B-Instruct, a powerful instruction-tuned language model. This model was used to assess the semantic clarity, descriptiveness, and relevance of each caption to ensure alignment with our image generation goals. Captions that met the predefined criteria for specificity, visual richness, and informativeness were retained.
Appendix C Another Advantage: Addressing Data Privacy
Our synthetic approach removes reliance on real-world personal data (e.g., user photos), safeguarding user privacy throughout the data generation process while maintaining model capability.
we compare the synthetic image and the original image in Figure 4. Synthetic data offers significant advantages in protecting data privacy. In Figure 4, synthetic images in (a) show vehicles and tickets without revealing real license plates and ticket information, ensuring privacy protection. In contrast, original images in (b) display actual license plates and ticket information, which can potentially lead to privacy issues.
Appendix D T-SNE visualize of our dataset
(a) Visualization of text embeddings.
(b) Visualization of image embeddings.
Figure 5. TSNE visualizations of synthetic and real datasets for text and image modalities.
In this section, we use t-distributed Stochastic Neighbor Embedding (t-SNE) to compare the distribution of our synthetic dataset with several real-world datasets. This comparison aims to evaluate the similarity in semantic and visual space, providing insights into the realism and utility of the generated data.
For the real datasets, we selected LLaVA-558K(Liu et al., 2023a), ShareCaptioner(Chen et al., 2023b), and CC12M(Singla et al., 2024), which are widely used for vision-language training. As our synthetic dataset, we used SynthVLM-100K, generated using the methods described in earlier sections. From each dataset, we randomly sampled 1k image-caption pairs for analysis.
We performed t-SNE visualization separately on image features and caption embeddings. Feature representations were extracted using a pre-trained vision-language model to ensure consistency and comparability across datasets.
As shown in Figure 5, the image and caption distributions of our synthetic dataset are closely aligned with those of the real datasets. This visual overlap indicates that the generated data captures similar semantic and visual characteristics as real-world data, supporting the authenticity and high quality of our generation pipeline.
Furthermore, the observed distributional similarity suggests that models trained on our synthetic data are likely to exhibit strong generalization and performance on real-world tasks. This supports the viability of using synthetic data to supplement or replace real data in various vision-language applications.
Appendix E More examples of our dataset
In this section, we present additional qualitative examples from our synthetic dataset, SynthVLM-100K, to further demonstrate the high quality and diversity of the generated image-caption pairs.
As illustrated in Figures 1 through 4, the samples cover a wide range of visual concepts and exhibit strong semantic alignment between images and captions. These examples highlight the capability of our data generation pipeline to produce visually coherent and semantically rich content across various domains.
Figure 6. Illustration of our SynthVLM
Figure 7. Illustration of our SynthVLM
Figure 8. Illustration of our SynthVLM
Figure 9. Illustration of our SynthVLM